OpenStudy (anonymous):

if we multiply two negetive numbers then we get positive number this why? what is the logic??

OpenStudy (sparrow2):

it is just like axiom,

Nnesha (nnesha):

bec \(\huge\color{green}{{\rm negative ~\times ~negative ~=positive}}\) that's why ;-) :D

OpenStudy (anonymous):

that is actually a very very good question the link between multiplication and addition is always the distributive property using it, you can show that \((-a)\times (-b)\) is the additive inverse of \(-a\times b\)

OpenStudy (sparrow2):

so why 2+2=4? in 10 numeral system

OpenStudy (sparrow2):

why not? :D

OpenStudy (anonymous):

Nnesha im satisfied with your ans.my question is that why this is happened??

OpenStudy (amistre64):

what is multiplication? what is it short cutting?

OpenStudy (amistre64):

lets say we want to add together 500 of the number 3 3+3+3+3+...+3 we can do it by addition, but its rather labor intensive. if we factor our a 3 (1+1+1+1+..+1)*3 = 500(3) multiplication is just faster addition

OpenStudy (anonymous):

Positive and negative are opposites of each other. In other words, negation of one thing results in other. i.e., if you negate a positive, it will become negative. On the similar terms, if you negate a negative, it will become positive. So, when u do, -1*-1, you are basically negating a negative number, which will ultimately return into a positive number.

OpenStudy (anonymous):

i know but what is the reason??

OpenStudy (anonymous):

- multiply - we get + any logic??

OpenStudy (amistre64):

say we want to add up 500 of the number -3 (-3) + (-3) + (-3) + ... +(-3) (1+1+1+1+..+1)*(-3) = 500(-3) --------------------------------- subtracting 500(-3) from itself should give us zero 500(-3) - 500(-3) = 500(-3-(-3)) = 0 --------------------------------- the 500 is superfluous -3 -(-3) = 0 we also know that -3+3 = 0 -3+3 = -3-(-3) , add 3 to each side 3 = -(-3)

OpenStudy (amistre64):

of course the number 3 is nothing special, and any real number suffices, say x x = -(-x)

OpenStudy (amistre64):

might be a couple of extra steps, but it still shows the logic